A soft-drink machine is regulated so that it discharges an average of 200 mm/cup. Of the amount of drink is normally distributed with a standard deviation equal to 15 mm, how many cups will probably overflow if 230 mm cups are used for the next 1000 drinks?


1
Expert's answer
2021-05-17T04:58:02-0400

"z= \frac{x-\mu}{\sigma} = \frac{230-200}{15} = 2"



The probability of overflow if 230 cups is used


P(X>230) = P(Z>2)

= 1- P(Z<2)

= 1- 0.9772= 0.0228




by using binomial property we get


E(X)= "\eta .p"


=1000 "\times 0.0228" = 22.8